A shopper in a supermarket pushes a loaded cart with a horizontal force of 100 N. The cart has a mass of 30 kg. How far will it move in 3.0 s if the shopper places his 30-N child in the cart before he begins to push it? The coefficient of kinetic friction is 0.2
1
Expert's answer
2020-12-21T11:01:13-0500

Let's first find acceleration from the Newtons Second Law of Motion:


"F_{appl}-F_{fr}=(m_{cart}+m_{child})a,""F_{appl}-\mu_k (m_{cart}+m_{child})g=(m_{cart}+m_{child})a,""a=\dfrac{F_{appl}-\mu_k (m_{cart}+m_{child})g}{(m_{cart}+m_{child})},""a=\dfrac{100\ N-0.2\cdot (30\ kg+\dfrac{30\ N}{9.8\ \dfrac{m}{s^2}})\cdot 9.8\ \dfrac{m}{s^2}}{(30\ kg+\dfrac{30\ N}{9.8\ \dfrac{m}{s^2}})}=1.1\ \dfrac{m}{s^2}."

Finally, we can find the distance traveled by the cart in 3 s from the kinematic equation:


"d=v_0t+\dfrac{1}{2}at^2=0+\dfrac{1}{2}\cdot 1.1\ \dfrac{m}{s^2}\cdot (3\ s)^2=4.95\ m."

Answer:

"d=4.95\ m."